You diagnosed 7 year old Emma with right otitis media. Emma weighs 81 lbs. You order amoxicillin 90 mg/kg a day divided every 8 hours for seven days. Amoxicillin chewable comes in 125 mg or 250 mg strength


What strength will you order?

How many chewable will you order?
Answer #1

Weight of Emma = 36.74 kg

Amoxycillin per kg in one day = 90 mg

Total Amoxycillin in one day = 3306.6 mg

Amoxycillin per 8 hourly = 1102.2mg

Strength should be 250 mg for less number of tablet

Tablet = 1102.2/ 250 4.5 approx

Total chewable in one day = 4.5*3 13.5 tablet

Chewable in 7 days,= 7*13.5 = 94 tablets